# Revisiting CNNs for Trajectory Similarity Learning

### v1

@@ Abstract.
In this paper, we argue that the common practice of treating trajectory as sequential data results in excessive attention to capturing long-term global dependency between two sequences. Instead, our investigation reveals the pivotal role of local similarity, prompting a revisit of simple CNNs for trajectory similarity learning. We introduce ConvTraj, incorporating both 1D and 2D convolutions to capture sequential and geo-distribution features of trajectories, respectively. In addition, we conduct a series of theoretical analyses to justify the effectiveness of ConvTraj. Experimental results on three real-world large-scale datasets demonstrate that ConvTraj achieves state-of-the-art accuracy in trajectory similarity search. Owing to the simple network structure of ConvTraj, the training and inference speed on the Porto dataset with 1.6 million trajectories are increased by at least $240$x and $2.16$x, respectively. The source code and dataset can be found at https://github.com/Proudc/ConvTraj.
@@ 1. Introduction
Method # Paras Inference time HR@1 HR@5 HR@10 HR@50 HR@1 HR@5 HR@10 HR@50
global attention 3.38M 3.58s $11.53_{\pm 0.01}$ $21.49_{\pm 0.93}$ $26.46_{\pm 0.62}$ $35.85_{\pm 0.73}$ $26.85_{\pm 3.06}$ $43.56_{\pm 0.03}$ $49.84_{\pm 0.29}$ $59.99_{\pm 0.47}$
local attention ($w=10$) 3.38M 3.58s $16.60_{\pm 0.01}$ $26.47_{\pm 0.55}$ $\textbf{31.89}_{\pm 1.48}$ $41.40_{\pm 1.92}$ $\textbf{31.03}_{\pm 5.42}$ $\textbf{47.42}_{\pm 0.50}$ $\textbf{52.73}_{\pm 0.38}$ $\textbf{62.88}_{\pm 0.89}$
local attention ($w=5$) 3.38M 3.58s $\textbf{17.07}_{\pm 0.34}$ $\textbf{26.75}_{\pm 0.32}$ $31.82_{\pm 1.52}$ $\textbf{41.53}_{\pm 1.93}$ $29.10_{\pm 0.29}$ $44.22_{\pm 0.64}$ $49.55_{\pm 0.53}$ $59.07_{\pm 0.30}$
1D CNN 0.17M 0.16s $\textbf{33.23}_{\pm 1.23}$ $\textbf{43.94}_{\pm 0.29}$ $\textbf{50.84}_{\pm 0.39}$ $\textbf{64.78}_{\pm 0.04}$ $30.90_{\pm 0.99}$ $46.66_{\pm 0.27}$ $\textbf{53.36}_{\pm 0.26}$ $\textbf{65.14}_{\pm 0.23}$
@@
These observations reveal the pivotal role of local similarity. Instead of adopting Transformer with masked local attention, we are interested in revisiting CNNs in the task of trajectory similarity learning. The reason is that CNNs can also well capture local similarity while offering the advantages of simplicity and smaller network structures. As shown in  Table 1, with only $5\%$ of the parameters, a simple 1D CNN can remarkably outperform Transformers. To further exploit the potential of CNNs, we present ConvTraj with two types of convolutions. We first use 1D convolution to capture the sequential features of trajectories. Then we represent the trajectory as a single-channel binary image and use 2D convolution to capture its geo-distribution. Finally, these features are fused as complementary clues to capture trajectory similarity. To justify the effectiveness of ConvTraj, we conduct a series of theoretical analyses. We prove that 1D convolution and 1D max-pooling can preserve effective distance bounds after embedding, and trajectories located in distant areas yield large distances via 2D convolution, all of which play an important role in trajectory similarity recognition.
We conducted extensive experiments to evaluate the performance of ConvTraj on three real-world datasets. Experimental results show that ConvTraj achieves state-of-the-art accuracy for similarity retrieval on four commonly used similarity measurements, including DFD, DTW, Hausdorff, and EDR. Furthermore, ConvTraj is at least $240$x faster in training speed and $2.16$x faster in inference speed, when compared with methods based on RNN and Transformer on the Porto dataset containing 1.6 million trajectories.
@@
  • Extensive experiments on three real-world large-scale datasets established the superiority of ConvTraj over state-of-the-art works in terms of accuracy and efficiency.
@@ 5. Theoretical Analysis
In this section, we will conduct some theoretical analysis from both 1D and 2D convolution to help justify why such a simple ConvTraj can perform well. We take the DFD, which is widely used for trajectory similarity (Zhang et al., 2022; Xie et al., 2017; Zhang et al., 2019; Tang et al., 2017), as an example for analysis. In summary, we found that: (1) For a randomly initialized kernel of 1D convolution, the DFD between two trajectories can still be maintained to a large extent. (2) After 1D max-pooling, the DFD value has almost no change. (3) Trajectories located in distant areas not only have a large DFD value but also have a large Euclidean distance through 2D convolution. Since 1D convolution essentially rotates and scales the sequence and 2D convolution captures the geo-distribution of trajectories, thus similar conclusions can be easily generalized to other measurements.
@@ 6.1. Experimental Setting
Datasets. We evaluate the performance of ConvTraj using three widely used real-world datasets: Geolife [^2] https://www.microsoft.com/en-us/research/publication/geolife-gps-trajectory-dataset-user-guide/, Porto [^3] https://www.kaggle.com/competitions/pkdd-15-predict-taxi-service-trajectory-i/data and TrajCL-Porto [^4] https://github.com/changyanchuan/TrajCL. For Geolife and Porto, we preprocess them using the method in (Yao et al., 2019), i.e. selecting trajectories in the central area of the city and removing items with less than 10 records. For TrajCL-Porto, it’s an open-source dataset of TrajCL(Chang et al., 2023a), we thus do not perform any processing. The properties of these datasets are shown in Table 2.
@@
Dataset Geolife Porto TrajCL-Porto
# Total Items $13,386$ $1,601,579$ $9000$
# Training Items $3000$ $3000$ $7000$
# Query Items $1000$ $500$ $2000$
# Candidate Items $9386$ $1,598,079$ $2000$
Avg-(# GPS Points) $437.80$ $48.91$ $49.72$
Min-(# GPS Points) $11$ $11$ $20$
Max-(# GPS Points) $7579$ $3836$ $200$
Lat and Lon Area (116.20, 116.50) (39.85, 40.07) (-8.73, -8.50) (41.10, 41.24) (-8.70, -8.52) (41.10, 41.208)
@@
Baselines. When we test on Geolife and Porto, we follow existing works (Yao et al., 2022; Chang et al., 2023c) and compare ConvTraj with six representative methods, including t2vec (Li et al., 2018) and TrjSR (Cao et al., 2021) based on self-supervised learning; NeuTraj (Yao et al., 2019), Traj2SimVec (Zhang et al., 2020), TrajGAT (Yao et al., 2022), and TrajCL (Chang et al., 2023a) based on supervised learning. For the self-supervised method, since its goal is not to approximate the existing measurements, we thus perform the following steps to handle it. We first randomly select a part of the trajectory for pre-training (We select 10000 trajectories for Geolife and 200000 for Porto. Since TrajCL also needs pre-training, we will use these data to pre-train TrajCL). Then we add an MLP encoder in the end and fine-tune it with the triplet selection method and loss function in  subsection 4.3. For those methods which have open-source code (Li et al., 2018; Chang et al., 2023a; Yao et al., 2019, 2022; Cao et al., 2021), we directly use their implementation. For others  (Zhang et al., 2020), we implement it based on the settings of its paper.
In addition, since many baselines have been evaluated on the TrajCL-Porto and the results have been reported in (Chang et al., 2023a), we will directly compare the results of ConvTraj on TrajCL-Porto with those of other baselines reported in (Chang et al., 2023a).
@@
Implementation Details. We set the MLP output dimension in 1D preprocessing to 16, and grid width $\delta=250m$ when generating binary images. For the Geolife dataset, the number of residual blocks for 1D convolution is $n=12(\lfloor\log_{2}7579\rfloor)$, Porto is $n=11(\lfloor\log_{2}3836\rfloor)$, and TrajCl-Porto is $n=7(\lfloor\log_{2}200\rfloor)$. During training, we set the batch size to 128, the learning rate to $0.001$, and the embedding dimension to 128. We evaluated four common trajectory similarity measurements, Hausdorff, DFD, DTW, and EDR on Geolife and Porto, and evaluated Hausdorff, DFD, EDwP (Ranu et al., 2015), and EDR on TrajCL-Porto. For each measurement on Geolife and Porto, we select three random seeds to repeat the experiment and report the average result and variance. All experiments are conducted on a machine equipped with 36 CPU cores (Intel Core i9-10980XE CPU with 3.00GHz), 256 GB RAM, and a GeForce RTX 3090Ti GPU.
@@ 6.2. Effectiveness
Table 3 and  Table 4 present an overview of the performance exhibited by different methods concerning the top-$k$ similarity search task on Geolife and Porto, we can observe that: (1) On both Geolife and Porto datasets, ConvTraj significantly outperforms all methods on all metrics. Taking the Hausdorff distance on the Geolife as an example, compared with the state-of-the-art baseline NeuTraj, ConvTraj exceeds by more than 11% in all metrics, with the largest improvement of 15.42% for HR@5 and the smallest improvement of 11.64% for HR@1. In addition, even for the Porto which contains 1.6 million trajectories, R10@50 has at least a 10.75% improvement on four measurements. This non-negligible improvement in performance is impressive given the fact that the sequence order features extracted by 1D convolution and the geographical distribution of the trajectory extracted by 2D convolution are both very beneficial to generating high-quality trajectory embedding representations. (2) The advantage of ConvTraj is evident in all measurements, which shows that ConvTraj is a general framework for different measurements. We can observe that no method can handle all measurements well. For example, NeuTraj performs best on the Hausdorff and DFD, while TrjSR and TrajCL have advantages on DTW and EDR respectively, which is also mutually verified with the results in  (Chang et al., 2023c). However, ConvTraj achieves state-of-the-art accuracy in all measurements. Compared to the state-of-the-art, ConvTraj achieves an average improvement of 10.22%, 8.02%, 7.59%, and 7.03% on all metrics of the Hausdorff, DFD, DTW, and EDR in Porto respectively.
@@
Table 5 presents the experimental results on TrajCL-Porto, we can observe that: (1) Similar to its performance on the Geolife and Porto, the ConvTraj method surpasses state-of-the-art in almost all metrics for four measurements. Compared with the state-of-the-art, Convtraj achieves improvements of 12%, 23%, 6.8%, and 14.2% on the HR@5 metrics of EDR, EDwP, Hausdorff, and DFD. (2) Even though both were tested on the Porto dataset, the performance gap between Table 3, Table 4 and Table 5 is very large. For example, the HR@5 of the TrajCL and ConvTraj in Table 5 on the DFD are 0.618 and 0.749 respectively, but in Table 4 they are 0.141 and 0.349 respectively. The reason is that the TrajCL-Porto dataset contains fewer trajectories. When performing the top-$k$ similarity search task, the TrajCL-Porto dataset only has 2000 candidate trajectories. However, the Porto used in Table 3 and Table 4 contains 1598079 candidate trajectories, which results in a more comprehensive result. (3) We also evaluate the performance of ConvTraj using only 1D convolution (ConvTraj-1D CNN) or 2D convolution (ConvTraj-2D CNN) on TrajCL-Porto, and we can observe that ConvTraj’s performance degrades after missing some features, but still has excellent performance.
@@ 6.3. Efficiency
We evaluate the efficiency of all baselines with open-source implementations and report the results in  Table 6. This includes network parameters, training time, and inference time. For methods that require pre-training, we also report their pre-training time. As illustrated, compared to existing RNN-based and Transformer-based methods, ConvTraj not only has fewer parameters (only 0.03M more than NeuTraj) but also has great advantages in training and inference speed. Taking the Porto with 1.6 million items as an example, compared with the most efficient Transformer-based model TrajCL, the training speed per epoch and the inference speed of ConvTraj are at least 243.80x and 12.87x faster respectively. Compared with the most efficient RNN-based model t2vec, the training speed per epoch and the inference speed of ConvTraj are at least 306.91x and 2.16x faster respectively. The reason for such a huge improvement is that compared to Transformer-based methods, ConvTraj has fewer parameters. Meanwhile, compared with RNN-based methods, although the parameters of ConvTraj are relatively large, the training and inference of ConvTraj are more efficient due to the inherent low parallelism of RNN. In addition, we also note that: (1) Compared with the CNN-based TrjSR, ConvTraj has no advantage in inference, but the training is faster because TrjSR requires pre-training on a large number of trajectories, and only uses fewer convolutional layers during inference, which also shows the superiority of CNN in terms of efficiency. (2) Although both t2vec and NeuTraj are based on RNN, and NeuTraj has fewer parameters, t2vec is more efficient. The reason is that NeuTraj needs to select more triplets during the training phase and compute spatial attention based on adjacent grids at each time step.

### v2

@@ Abstract.
In this paper, we argue that the common practice of treating trajectory as sequential data results in excessive attention to capturing long-term global dependency between two sequences. Instead, our investigation reveals the pivotal role of local similarity, prompting a revisit of simple CNNs for trajectory similarity learning. We introduce ConvTraj, incorporating both 1D and 2D convolutions to capture sequential and geo-distribution features of trajectories, respectively. In addition, we conduct a series of theoretical analyses to justify the effectiveness of ConvTraj. Experimental results on four real-world large-scale datasets demonstrate that ConvTraj achieves state-of-the-art accuracy in trajectory similarity search. Owing to the simple network structure of ConvTraj, the training and inference speed on the Porto dataset with 1.6 million trajectories are increased by at least $240$x and $2.16$x, respectively. The source code and dataset can be found at https://github.com/Proudc/ConvTraj.
@@ 1. Introduction
Method # Paras (Train time Per Epoch) * # Epochs Inference time HR@1 HR@5 HR@10 HR@50 HR@1 HR@5 HR@10 HR@50
global attention 3.38M 17.28s * 1000 3.58s $22.10$ $32.58$ $39.11$ $50.11$ $29.40$ $46.22$ $54.60$ $63.41$
local attention ($w=10$) 3.38M 17.28s * 1000 3.58s 23.20 36.74 42.80 54.70 $31.50$ 48.60 54.92 65.06
local attention ($w=5$) 3.38M 17.28s * 1000 3.58s $21.80$ $35.40$ $41.83$ $54.42$ 33.60 $46.72$ $52.34$ $63.03$
1D CNN 0.17M 1.03s * 200 0.16s 33.23 43.94 50.84 64.78 $30.90$ $46.66$ $53.36$ 65.14
@@
These observations reveal the pivotal role of local similarity. Instead of adopting Transformer with masked local attention, we are interested in revisiting CNNs in the task of trajectory similarity learning. The reason is that CNNs can also well capture local similarity while offering the advantages of simplicity. As shown in  Table 1, with only 5% of the parameters, a simple 1D CNN can remarkably outperform vanilla Transformers after convergence on the DFD. Although slightly lower than local attention on the DTW, 1D CNN has great advantages in efficiency. To further exploit the potential of CNNs, we present ConvTraj with two types of convolutions. We first use 1D convolution to capture the sequential features of trajectories. Then we represent the trajectory as a single-channel binary image and use 2D convolution to capture its geo-distribution. Finally, these features are fused as complementary clues to capture trajectory similarity. To justify the effectiveness of ConvTraj, we conduct a series of theoretical analyses. We prove that 1D convolution and 1D max-pooling can preserve effective distance bounds after embedding, and trajectories located in distant areas yield large distances via 2D convolution, all of which play an important role in trajectory similarity recognition.
We conducted extensive experiments to evaluate the performance of ConvTraj on four real-world datasets. Experimental results show that ConvTraj achieves state-of-the-art accuracy for similarity retrieval on four commonly used similarity measurements, including DFD, DTW, Hausdorff, and EDR. Furthermore, ConvTraj is at least $240$x faster in training speed and $2.16$x faster in inference speed, when compared with methods based on RNN and Transformer on the Porto dataset containing 1.6 million trajectories.
@@
  • Extensive experiments on four real-world large-scale datasets established the superiority of ConvTraj over state-of-the-art works in terms of accuracy and efficiency.
@@ 5. Theoretical Analysis
In this section, we will conduct some theoretical analysis from both 1D and 2D convolution to help justify why such a simple ConvTraj can perform well. We take the DFD, which is widely used for trajectory similarity (Zhang et al., 2022; Xie et al., 2017; Zhang et al., 2019; Tang et al., 2017), as an example for analysis. In summary, we found that: (1) For a randomly initialized kernel of 1D convolution, the DFD between two trajectories can still be maintained to a large extent. (2) After 1D max-pooling, the DFD value has almost no change. (3) Trajectories located in distant areas not only have a large DFD value but also have a large Euclidean distance through 2D convolution. Since 1D convolution essentially rotates and scales the sequence and 2D convolution captures the geo-distribution of trajectories, thus similar conclusions can be easily generalized to other measurements. Basically, the analysis shows that 1D convolution and max-pooling can preserve the bounds for trajectory similarity learning, while 2D convolution can help capture the geo-distribution. This implies CNNs are a good choice in scenarios where trajectories need to be reduced in dimension or geo-distribution is required. This does not mean that RNNs or Transformers lack it, it is just difficult to analyze.
@@ 6.1. Experimental Setting
Datasets. We evaluate the performance of ConvTraj using four widely used real-world datasets: Geolife [^2] https://www.microsoft.com/en-us/research/publication/geolife-gps-trajectory-dataset-user-guide/, Porto [^3] https://www.kaggle.com/competitions/pkdd-15-predict-taxi-service-trajectory-i/data, Chengdu and TrajCL-Porto [^4] https://github.com/changyanchuan/TrajCL. For Geolife and Porto, we preprocess them using the method in (Yao et al., 2019), i.e. selecting trajectories in the central area of the city and removing items with less than 10 records. For TrajCL-Porto, it’s an open-source dataset of TrajCL(Chang et al., 2023a), we thus do not perform any processing. For Chengdu, we randomly selected 5000 trajectories from this dataset. The properties of these datasets are shown in Table 2.
@@
Dataset Geolife Porto Chengdu TrajCL-Porto
# Total Items $13386$ $1601579$ $5000$ $9000$
# Training Items $3000$ $3000$ $1000$ $7000$
# Query Items $1000$ $500$ $1000$ $2000$
# Candidate Items $9386$ $1598079$ $3000$ $2000$
Avg-(# Points) $437.80$ $48.91$ $228.44$ $49.72$
Min-(# Points) $11$ $11$ $29$ $20$
Max-(# Points) $7579$ $3836$ $1575$ $200$
Lat-Lon Area (116.20, 116.50) (39.85, 40.07) (-8.73, -8.50) (41.10, 41.24) (104.04, 114.10) (30.65, 30.73) (-8.70, -8.52) (41.10, 41.208)
@@
Table 5. Embedding Results On Chengdu dataset
Chengdu
Hausdorff DFD DTW EDR
Model HR@5 HR@10 R10@50 HR@5 HR@10 R10@50 HR@5 HR@10 R10@50 HR@5 HR@10 R10@50
t2vec $16.16$ $16.52$ $30.69$ $21.34$ $22.34$ $44.06$ $21.00$ $22.68$ $46.03$ $14.18$ $16.37$ $37.41$
TrjSR $10.44$ $12.09$ $25.44$ $11.08$ $12.77$ $25.96$ $16.42$ $18.32$ $36.63$ $\underline{18.80}$ $19.78$ $40.02$
TrajCL $\underline{22.14}$ $\underline{27.04}$ $\underline{61.93}$ $19.00$ $21.42$ $51.79$ $23.48$ $27.20$ $59.89$ $17.40$ $20.53$ $\underline{48.58}$
NeuTraj $21.82$ $23.27$ $39.09$ $\underline{32.28}$ $\underline{34.70}$ $\underline{49.51}$ $\underline{28.40}$ $\underline{29.33}$ $46.90$ $10.44$ $11.54$ $24.65$
Traj2SimVec $18.34$ $20.17$ $43.67$ $25.16$ $28.33$ $42.78$ $18.66$ $19.37$ $40.69$ $6.93$ $8.31$ $19.18$
TrajGAT $19.98$ $25.26$ $57.57$ $16.24$ $19.11$ $49.18$ $23.96$ $28.32$ $\underline{65.41}$ $17.94$ $\underline{20.55}$ $48.40$
ConvTraj 36.26 42.78 76.67 53.34 58.18 93.14 34.90 40.40 76.23 21.50 25.34 55.21
Gap with SOTA $+14.12$ $+15.74$ $+19.10$ $+21.06$ $+23.48$ $+41.45$ $+6.50$ $+11.07$ $+10.82$ $+2.70$ $+4.81$ $+6.63$
@@
Baselines. When we test on Geolife, Porto, and Chengdu, we follow existing works (Yao et al., 2022; Chang et al., 2023c) and compare ConvTraj with six representative methods, including t2vec (Li et al., 2018) and TrjSR (Cao et al., 2021) based on self-supervised learning; NeuTraj (Yao et al., 2019), Traj2SimVec (Zhang et al., 2020), TrajGAT (Yao et al., 2022), and TrajCL (Chang et al., 2023a) based on supervised learning. For the self-supervised method, since its goal is not to approximate the existing measurements, we thus perform the following steps to handle it. We first randomly select a part of the trajectory for pre-training (We select 10000 trajectories for Geolife and 200000 for Porto. Since TrajCL also needs pre-training, we will use these data to pre-train TrajCL). Then we add an MLP encoder in the end and fine-tune it with the triplet selection method and loss function in  subsection 4.3. For those methods which have open-source code (Li et al., 2018; Chang et al., 2023a; Yao et al., 2019, 2022; Cao et al., 2021), we directly use their implementation. For others  (Zhang et al., 2020), we implement it based on the settings of its paper. In addition, since many baselines have been evaluated on the TrajCL-Porto and the results have been reported in (Chang et al., 2023a), we will directly compare our results with those of other baselines reported in (Chang et al., 2023a).
@@
Implementation Details. We set the MLP output dimension in 1D preprocessing to 16. Intuitively, as the grid width $\delta$ decreases, ConvTraj will perform better, but the training cost of the model will also increase significantly. We thus set $\delta$ as 250 meters when generating binary images. For the Geolife dataset, the number of residual blocks for 1D convolution is $n=12(\lfloor\log_{2}7579\rfloor)$, Porto is $n=11(\lfloor\log_{2}3836\rfloor)$, and TrajCl-Porto is $n=7(\lfloor\log_{2}200\rfloor)$. During training, we set the batch size to 128, the learning rate to $0.001$, and the embedding dimension to 128. We evaluated four common trajectory similarity measurements, Hausdorff, DFD, DTW, and EDR on Geolife and Porto, and evaluated Hausdorff, DFD, EDwP (Ranu et al., 2015), and EDR on TrajCL-Porto. For each measurement on Geolife and Porto, we select three random seeds to repeat the experiment and report the average result and variance. All experiments are conducted on a machine equipped with 36 CPU cores (Intel Core i9-10980XE CPU with 3.00GHz), 256 GB RAM, and a GeForce RTX 3090Ti GPU.
@@ 6.2. Effectiveness
Table 3,  Table 4, and Table 5 present an overview of the performance exhibited by different methods concerning the top-$k$ similarity search task on Geolife, Porto, and Chengdu, we can observe that: (1) On all datasets, ConvTraj significantly outperforms all methods on all metrics. Taking the Hausdorff distance on the Geolife as an example, compared with the state-of-the-art baseline NeuTraj, ConvTraj exceeds by more than 11% in all metrics, with the largest improvement of 15.42% for HR@5 and the smallest improvement of 11.64% for HR@1. In addition, even for the Porto which contains 1.6 million trajectories, R10@50 has at least a 10.75% improvement on four measurements. This non-negligible improvement in performance is impressive given the fact that the sequence order features extracted by 1D convolution and the geographical distribution of the trajectory extracted by 2D convolution are both very beneficial to generating high-quality trajectory embedding representations. (2) The advantage of ConvTraj is evident in all measurements, which shows that ConvTraj is a general framework for different measurements. We can observe that no method can handle all measurements well. For example, NeuTraj performs best on the Hausdorff and DFD, while TrjSR and TrajCL have advantages on DTW and EDR respectively, which is also mutually verified with the results in  (Chang et al., 2023c). However, ConvTraj achieves state-of-the-art accuracy in all measurements. Compared to the state-of-the-art, ConvTraj achieves an average improvement of 10.22%, 8.02%, 7.59%, and 7.03% on all metrics of the Hausdorff, DFD, DTW, and EDR in Porto respectively. (3) We also noticed that compared with the results on the Geolife and Porto datasets, the TrajGAT method performed better on the Chengdu dataset. This may be because the longitude and latitude of the Chengdu dataset cover a larger area, so the quadtree-based modeling method of the TrajGAT is more effective.
@@
Table 6 presents the experimental results on TrajCL-Porto, we can observe that: (1) Similar to its performance on the Geolife and Porto, the ConvTraj method surpasses state-of-the-art in almost all metrics for four measurements. Compared with the state-of-the-art, Convtraj achieves improvements of 12%, 23%, 6.8%, and 14.2% on the HR@5 metrics of EDR, EDwP, Hausdorff, and DFD. (2) Even though both were tested on the Porto dataset, the performance gap between Table 4 and Table 6 is very large. For example, the HR@5 of the TrajCL and ConvTraj in Table 6 on the DFD are 0.618 and 0.760 respectively, but in Table 4 they are 0.141 and 0.349 respectively. The reason is that the TrajCL-Porto dataset contains fewer trajectories. When performing the top-$k$ similarity search task, the TrajCL-Porto dataset only has 2000 candidate trajectories. However, the Porto used in Table 3 and Table 4 contains 1598079 candidate trajectories, which results in a more comprehensive result. (3) We also evaluate the performance of ConvTraj using only 1D convolution (ConvTraj-1D CNN) or 2D convolution (ConvTraj-2D CNN) on TrajCL-Porto, and we can observe that ConvTraj’s performance degrades after missing some features, but still has excellent performance.
@@ 6.3. Efficiency
We evaluate the efficiency of all baselines with open-source implementations on Geolife and Porto, and report the results in  Table 7. This includes network parameters, training time, and inference time. For methods that require pre-training, we also report their pre-training time. As illustrated, compared to existing RNN-based and Transformer-based methods, ConvTraj not only has fewer parameters (only 0.03M more than NeuTraj) but also has great advantages in training and inference speed. Taking the Porto with 1.6 million items as an example, compared with the most efficient Transformer-based model TrajCL, the training speed per epoch and the inference speed of ConvTraj are at least 243.80x and 12.87x faster respectively. Compared with the most efficient RNN-based model t2vec, the training speed per epoch and the inference speed of ConvTraj are at least 306.91x and 2.16x faster respectively. The reason for such a huge improvement is that compared to Transformer-based methods, ConvTraj has fewer parameters. Meanwhile, compared with RNN-based methods, although the parameters of ConvTraj are relatively large, the training and inference of ConvTraj are more efficient due to the inherent low parallelism of RNN. In addition, we also note that: (1) Compared with the CNN-based TrjSR, ConvTraj has no advantage in inference, but the training is faster because TrjSR requires pre-training on a large number of trajectories, and only uses fewer convolutional layers during inference, which also shows the superiority of CNN in terms of efficiency. (2) Although both t2vec and NeuTraj are based on RNN, and NeuTraj has fewer parameters, t2vec is more efficient. The reason is that NeuTraj needs to select more triplets during the training phase and compute spatial attention based on adjacent grids at each time step.